OpenStudy (czarluc):

Exponential growth the formulas for growth y=a(1+r)^x and y=a(e^kx) where y=result a=starting amt. r=rate x=time elapsed k=constant any explanation when to use one of these???

OpenStudy (anonymous):

here is an example

OpenStudy (anonymous):

Example: 2 months ago you had 3 mice, you now have 18. Mice Assuming the growth continues like that What is the "k" value? How many mice 2 Months from now? How many mice 1 Year from now? Start with the formula: y(t) = a × ekt We know a=3 mice, t=2 months, and right now y(2)=18 mice: 18 = 3 × e2k Now some algebra to solve for k: Divide both sides by 3: 6 = e2k Take the natural logarithm of both sides: ln(6) = ln(e2k) ln(ex)=x, so: ln(6) = 2k Rearrange: k = ln(6)/2 Now, we want to know the population in 2 more months (at t=4 months), and in 1 year from now (t=14 months): y(4) = 3 e(ln(6)/2)×4 = 108 y(14) = 3 e(ln(6)/2)×14 = 839,808 That's a lot of mice! I hope you will be feeding them properly.

OpenStudy (anonymous):